# A preorder on the set of links defined via orbifolds

### v1

@@ 7. Symmetric unions of knots
8. Relation with other preorders on knots in $S^{3}$ and questions
In this section, we study the relationship of the $\pi$-domination with other preorders defined on the set of knots in $S^{3}$.
8.1. Relation with other preorders on knots in $S^{3}$
A classical domination relation between knots in $S^{3}$ is given as follows: a knot $K_{1}$ dominates another knot $K_{2}$ ($K_{1}\geqslant K_{2}$) if there is an epimorphism $\varphi\colon G(K_{1})\twoheadrightarrow G(K_{2})$. It is clear that any knot $K$ dominates the unknot $U$, and a connected sum $K_{1}\sharp K_{2}$ of knots does each summand.
This relation $\geqslant$ provides a partial order on the set of prime knots in $S^{3}$ up to mirror image. The reflexivity and transitivity of $\geqslant$ are clear. The anti-symmetry follows from the fact that knot groups are hopfian and that prime knots are determined by their groups. This is no longer true for a connected sum of knots. Moreover, this partial order cannot be extended to the setting of links since there are infinitely many distinct hyperbolic links with isomorphic link groups. For example, the twisted Whitehead links $W_{n}$ with $n\geq 0$ full twists are not equivalent, but the groups $G(W_{n})$ are isomorphic.
This domination relation can be refined to define for each integer $k\geq 1$ a $k$-domination relation $\geqslant_{k}$ in the following way: define the meridional complexity of an element $\gamma$ in a knot group $G(K)$ as the minimal number of distinct meridians in $G(K)$ needed to write $\gamma$ as a product of meridians. We recall that a meridian in $G(K)$ is any element which is conjugate to an element represented by a meridional simple closed curve on $\partial E(K)$. Then a knot $K_{1}$ $k$-dominates another knot $K_{2}$ ($K_{1}\geqslant_{k}K_{2}$) if there is an epimorphism $\varphi\colon G(K_{1})\twoheadrightarrow G(K_{2})$ which sends a meridian of $G(K_{1})$ to an element of meridional complexity being at most $k$ in $G(K_{2})$. The meridional complexity of the image of a meridian of $G(K_{1})$ does not depend on the choice of a meridian since all meridians are conjugate in $G(K_{1})$. From the definition, it is clear that $K_{1}\geqslant_{k}K_{2}$ implies $K_{1}\geqslant_{k^{\prime}}K_{2}$ for any $k^{\prime}\geq k$. Moreover, $K_{1}\geqslant K_{2}$ implies $K_{1}\geqslant_{k}K_{2}$ with $k\leq m(K_{2})$, where $m(K_{2})$ is the minimal number of meridians needed to generate $G(K_{2})$ which is called the meridional rank of $K_{2}$.
@@ 8.1. Relation with other preorders on knots in $S^{3}$
When $k=1$, the $1$-domination relation corresponds to epimorphisms between knot groups which send a meridian to a meridian or its inverse.
@@
Let us assume that $G(K_{1})\geqslant_{1}G(K_{2})$, then there is an epimorphism $\varphi\colon G(K_{1})\twoheadrightarrow G(K_{2})$ such that for a meridian $\mu_{1}\in G(K_{1})$ we have $\varphi(\mu_{1})=\mu_{2}^{n}$ for some meridian $\mu_{2}\in G(K_{2})$ and $n\in\mathbb{Z}\setminus\{0\}$. The epimorphism $\varphi$ induces an epimorphism
$\varphi_{\sharp}\colon G(K_{1})^{\mathrm{ab}}=\langle[\mu_{1}]\rangle\cong\mathbb{Z}\twoheadrightarrow G(K_{2})^{\mathrm{ab}}=\langle[\mu_{2}]\rangle\cong\mathbb{Z},$
where $\mu_{i}]\in G(K_{i})^{\mathrm{ab}}$ is the image of the meridian $\mu_{i}$ for $i=1,2$. The image $\varphi_{\sharp}([\mu_{1}])=n[\mu_{2}]$ generates $G(K_{2})^{\mathrm{ab}}=\langle[\mu_{K_{2}}]\rangle$, hence $n=\pm 1$. ∎
The the $1$-domination relation induces a partial order $\geqslant_{1}$ on the set of all knots in $S^{3}$ up to mirror image because knots are determined by their knot groups together with meridians. This partial order can be extended to the setting of links in $S^{3}$ since a link $L\subset S^{3}$ is determined by $G(L)$ together with a meridian for each component of $L$.
By definition, $K_{1}\geqslant_{1}K_{2}$ implies both $K_{1}\succeq K_{2}$ and $K_{1}\geqslant K_{2}$. However, the converse does not hold. In fact, while $K_{T}\succeq 3_{1}$ and $K_{T}\geqslant 3_{1}$ for the knot $K_{T}$ in [19, Section 2.1], we have $K_{T}\not\geqslant_{1}3_{1}$. By [19, Theorem 4.5], there are infinitely many such examples.
Moreover, $K_{1}\succeq K_{2}$ does not imply $K_{1}\geqslant K_{2}$. For example, $6_{1}\succeq 3_{1}$ since $6_{1}$ is a symmetric union with partial knot $3_{1}$, but $6_{1}\not\geqslant 3_{1}$ since the Alexander polynomial of $6_{1}$ is not divisible by that of $3_{1}$. It is still an open question whether $K_{1}\geqslant K_{2}$ implies $K_{1}\succeq K_{2}$. The following result shows that $K_{1}\geqslant_{2}K_{2}$ implies $K_{1}\succeq K_{2}$.
Proposition 8.2.
Let $\varphi\colon G(K_{1})\twoheadrightarrow G(K_{2})$ be an epimorphism between two knot groups. If the meridional complexity of the image of a meridian of $G(K_{1})$ is at most $2$, then $\varphi$ induces an epimorphism $G^{\mathrm{orb}}(K_{1})\twoheadrightarrow G^{\mathrm{orb}}(K_{2})$ between their $\pi$-orbifold groups.
@@
Let $\varphi\colon G(K_{1})\to G(K_{2})$ be an epimorphism such that for a meridian $\mu\in G(K_{1})$, $\varphi(\mu)=x^{p_{1}}y^{q_{1}}\cdots x^{p_{r}}y^{q_{r}}$ for two meridians $x,y$ in $G(K_{2})$ and some $p_{i},q_{i}\in\mathbb{Z}$. The epimorphism $\varphi$ induces an isomorphism
$\varphi_{\ast}\colon G(K_{1})^{\mathrm{ab}}=\langle[\mu]\rangle\cong\mathbb{Z}\to G(K_{2})^{\mathrm{ab}}=\langle[x]\rangle\cong\mathbb{Z}$
with $\varphi_{\ast}([\mu])=\big(\sum_{i=1}^{r}(p_{i}+q_{i})\big)[x]$. Therefore, $\sum_{i=1}^{r}(p_{i}+q_{i})=\pm 1$. It follows that $\varphi$ induces an epimorphism $\hat{\varphi}\colon G(K_{1})\twoheadrightarrow G(K_{2})\twoheadrightarrow G^{\mathrm{orb}}(K_{2})$ such that $\hat{\varphi}(\mu)$ is a product of an odd number of images of the meridians $x$ and $y$ in $G^{\mathrm{orb}}(K_{2})$. Hence $\hat{\varphi}(\mu)=\bar{x}\bar{y}\cdots\bar{x}$ or $\hat{\varphi}(\mu)=\bar{y}\bar{x}\cdots\bar{y}$. It is easy to verify that in both cases $\hat{\varphi}(\mu^{2})=1$ in $G^{\mathrm{orb}}(K_{2})$, and hence $\hat{\varphi}$ induces an epimorphism $G^{\mathrm{orb}}(K_{1})\twoheadrightarrow G^{\mathrm{orb}}(K_{2})$. ∎
The following corollary is straightforward since $2$-bridge knot groups are generated by two meridians.
Corollary 8.3.
If $K_{2}$ is a $2$-bridge knot, then every epimorphism $G(K_{1})\twoheadrightarrow G(K_{2})$ induces an epimorphism $G^{\mathrm{orb}}(K_{1})\twoheadrightarrow G^{\mathrm{orb}}(K_{2})$. In particular, $K_{1}\geqslant K_{2}$ implies $K_{1}\succeq K_{2}$.
It is worth noting that a $2$-domination $K_{1}\geqslant_{2}K_{2}$ does not imply a $1$-domination $K_{1}\geqslant_{1}K_{2}$. As mentioned above, we have $K_{T}\geqslant_{2}3_{1}$ since $3_{1}$ is a 2-bridge knot, but $K_{T}\not\geqslant_{1}3_{1}$.
There exist plenty of epimorphisms between knot groups which do not preserve meridians (see [19]). If $G(K)=\langle\!\langle\alpha\rangle\!\rangle$, is normally generated by an element $\alpha$ which is not an automorphic image of a meridian, then there is a knot $K^{\prime}$ and an epimorphism $\varphi\colon G(K^{\prime})\twoheadrightarrow G(K)$ such that $\varphi(\mu_{K^{\prime}})=\alpha$ (see [28] and [37]). Such an element $\alpha\in G(K)$ is called a pseudo-meridian for the knot group $G(K)$. It is conjectured that knot groups contain infinitely many inequivalent pseudo-meridians. This is true for $2$-bridge knots, torus knots, or hyperbolic knots with unknotting number one (see [60]).
8.2. Questions on bridge numbers

### v2

@@ 7.1. Symmetric unions of knots
A knot whose $\pi$-orbifold group is dihedral or $\mathbb{Z}/2\mathbb{Z}$ is a $2$-bridge knot or the unknot. Hence, the next corollary follows from Proposition 7.6 and the definition of $\pi$-minimality.
Corollary 7.9.
If a symmetric union $K$ with partial knot $K_{D}$ is $\pi$-minimal, then $K_{D}$ has at most two bridges.
Therefore, to prove Theorem 1.12, it remains to show that the condition on the rank of $H_{1}(\Sigma_{2}(K);\mathbb{Z})$ rules out the possibility of a partial knot $K_{D}$ with $\leq 2$ bridges. This follows from the next proposition since $H_{1}(\Sigma_{2}(K);\mathbb{Z})$ is cyclic or trivial for a knot with $\leq 2$ bridges.
Recall that a finite abelian group $A$ is isomorphic to a direct sum of finite cyclic groups: $G\cong\mathbb{Z}/d_{1}\mathbb{Z}\oplus\cdots\oplus\mathbb{Z}/d_{r}\mathbb{Z}$, where $r=\operatorname{rank}(A)$ and $d_{i}$ $(\geq 2)$ divides $d_{i+1}$ for $1\leq i\leq r-1$. The invariant factors $\{d_{1},\ldots,d_{r}\}$ are unique. Let $p$ be a prime number dividing $d_{1}$ (hence, $p\mid d_{i}$ for $2\leq i\leq r$) and let $\mathbb{F}_{p}$ denote the finite field of order $p$. Then, $\dim_{\mathbb{F}_{p}}(A\otimes\mathbb{F}_{p})=r$.
If $K$ admits a symmetric union presentation with partial knot $K_{D}$, then it is proved in [19, Theorem 1.3] that
$\mathrm{col}_{p}(K_{D})\leq\mathrm{col}_{p}(K)\leq\frac{\mathrm{col}_{p}(K_{D})^{2}}{p},$
where $p$ is an odd prime and $\mathrm{col}_{p}(K)$ denotes the number of $p$-colorings for any diagram of $K$. Since $\mathrm{col}_{p}(K)=p^{1+\dim_{p}H_{1}(\Sigma_{2}(K);\mathbb{F}_{p})}$, this result implies some inequalities between the mod $p$ first Betti numbers of a symmetric union and of its partial knot. This is the content of the next proposition for which we give a direct proof (which is also valid for $p=2$).
Proposition 7.10.
If $K$ admits a symmetric union presentation with partial knot $K_{D}$, then
$\operatorname{rank}H_{1}(\Sigma_{2}(K_{D});\mathbb{Z})\leq\operatorname{rank}H_{1}(\Sigma_{2}(K);\mathbb{Z})\leq 2\operatorname{rank}H_{1}(\Sigma_{2}(K_{D});\mathbb{Z}).$
Proof.
The epimorphism $\varphi_{D}^{\mathrm{orb}}\colon G^{\mathrm{orb}}(K)\twoheadrightarrow G^{\mathrm{orb}}(K_{D})$ induces an epimorphism $\tilde{\varphi}_{D}\colon\pi_{1}(\Sigma_{2}(K))\twoheadrightarrow\pi_{1}(\Sigma_{2}(K_{D}))$. Therefore, one gets the first inequality in the statement.
The proof of the second inequality uses the fact that the Goeritz matrix of a diagram of a knot $K$ is a presentation matrix for $H_{1}(\Sigma_{2}(K);\mathbb{Z})$. In [42, Section 3.2], it is proved that the Goeritz matrix $G_{K}$ for a symmetric union diagram of $K$ with partial knot $K_{D}$ is equivalent to $\begin{pmatrix}G_{D}&\ast\\ O&-G_{D}\end{pmatrix}$ by a combination of elementary row and column operations, where $G_{D}$ and $G_{D}$ are the Goeritz matrices for the knot diagram $D$ and $D^{*}$, respectively. Then, for any field $\mathbb{F}$, one gets $\operatorname{rank}_{\mathbb{F}}G_{K}\geq 2\operatorname{rank}_{\mathbb{F}}G_{D}$, and thus
$\dim_{\mathbb{F}}H_{1}(\Sigma_{2}(K);\mathbb{F})\leq 2\dim_{\mathbb{F}}H_{1}(\Sigma_{2}(K_{D});\mathbb{F}).$
Letting $\mathbb{F}=\mathbb{F}_{p}$ for a suitable prime $p$, we obtain the second inequality $\operatorname{rank}H_{1}(\Sigma_{2}(K);\mathbb{Z})\leq 2\operatorname{rank}H_{1}(\Sigma_{2}(K_{D});\mathbb{Z})$. ∎
Corollary 7.11.
Let $K$ be a knot such that $\operatorname{rank}H_{1}(\Sigma_{2}(K);\mathbb{Z})\geq 3$. If $K$ admits a symmetric union presentation with partial knot $K_{D}$, then $H_{1}(\Sigma_{2}(K_{D});\mathbb{Z})$ cannot be cyclic.
We call a knot $K\subset S^{3}$ almost $\pi$-minimal if $K\succeq K^{\prime}$ implies that $G^{\mathrm{orb}}(K^{\prime})$ is either finite or isomorphic to $G^{\mathrm{orb}}(K)$. If $G^{\mathrm{orb}}(K^{\prime})$ is finite, then $H_{1}(\Sigma_{2}(K^{\prime});\mathbb{Z})$ is cyclic, see [54, Theorem 2(iii) and (v) in Section 6.2]. Therefore, Proposition 7.6 and Corollary 7.11 imply the following improvement of Theorem 1.12.
Corollary 7.12.
Let $K$ be an almost $\pi$-minimal knot. If $\operatorname{rank}H_{1}(\Sigma_{2}(K);\mathbb{Z})\geq 3$, then $K$ cannot admit a symmetric union presentation.
For example, the proof of [5, Proposition 5.1] shows that the ribbon knot $11a_{201}$ is almost $\pi$-minimal, but $H_{1}(\Sigma_{2}(11a_{201});\mathbb{Z})\cong\mathbb{Z}/81\mathbb{Z}$ is cyclic.
It is still open whether the ribbon Montesinos knot $11a_{103}=K(\frac{2}{3},\frac{1}{3},\frac{2}{7})$ admits a symmetric union presentation or not. The next proposition shows that this knot is $\pi$-minimal. Hence, by Corollary 7.9, the only possible partial knots for a symmetric union presentation of the knot $11a_{103}$ are two $2$-bridge knots $6_{1}$ or $9_{1}$ because $\det(11a_{103})=81$.
Proposition 7.13.
The ribbon Montesinos knot $11a_{103}=K(\frac{2}{3},\frac{1}{3},\frac{2}{7})$ is $\pi$-minimal.
Proof.
Let assume that the knot $11a_{103}$ $\pi$-dominates a knot $K$. Since $11a_{103}$ is a Montesinos knot with three rational tangles, by Corollary 5.4, $K$ is either a knot with $\leq 2$ bridges or a Montesinos knot with three rational tangles.
Suppose that $K$ is a Montesinos knot $K(\frac{\beta_{1}}{\alpha_{1}},\frac{\beta_{2}}{\alpha_{2}},\frac{\beta_{3}}{\alpha_{3}})$. Then the epimorphism $\varphi\colon G^{\mathrm{orb}}(11a_{103})\twoheadrightarrow G^{\mathrm{orb}}(K)$ induces an epimorphism $\tilde{\varphi}\colon\pi_{1}(\Sigma_{2}(11a_{103}))\twoheadrightarrow\pi_{1}(\Sigma_{2}(K))$. Here, the $2$-fold cover $\Sigma_{2}(K)=V(0;e_{0};\frac{\beta_{1}}{\alpha_{1}},\frac{\beta_{2}}{\alpha_{2}},\frac{\beta_{3}}{\alpha_{3}})$ is a Seifert fibered manifold with base $S^{2}(\alpha_{1},\alpha_{2},\alpha_{3})$ and rational Euler number $e_{0}=\sum_{i=1}^{3}\frac{\beta_{i}}{\alpha_{i}}=\frac{|H_{1}(\Sigma_{2}(K);\mathbb{Z})|}{\alpha_{1}\alpha_{2}\alpha_{3}}$. In the same way, $\Sigma_{2}(11a_{103})=V(0;\frac{9}{7};\frac{2}{3},\frac{1}{3},\frac{2}{7})$ is Seifert fibered with base $S^{2}(3,3,7)$ and rational Euler number $\frac{9}{7}=\frac{|H_{1}(\Sigma_{2}(K);\mathbb{Z})|}{63}$, and hence $|H_{1}(\Sigma_{2}(K);\mathbb{Z})|=81$.
The epimorphism $\tilde{\varphi}$ induces an epimorphism $\bar{\varphi}\colon T(3,3,7)\twoheadrightarrow T(\alpha_{1},\alpha_{2},\alpha_{3})$ through taking the quotients by the centers. The presentation of the triangle group $T(3,3,7)=\langle x,y,z\mid x^{3}=y^{3}=z^{7}=xyz=1\rangle$ shows that each image $\bar{\varphi}(x)$, $\bar{\varphi}(y)$, and $\bar{\varphi}(z)$ is non-trivial, otherwise the image $\bar{\varphi}(T(3,3,7))$ would be either trivial or $\mathbb{Z}/3\mathbb{Z}$. Hence, $\bar{\varphi}(T(3,3,7))=T(\alpha_{1},\alpha_{2},\alpha_{3})$ must contain elements of orders $3$ and $7$. We distinguish two cases according to whether $\frac{1}{\alpha_{1}}+\frac{1}{\alpha_{2}}+\frac{1}{\alpha_{3}}\geq 1$ or not.
If $\frac{1}{\alpha_{1}}+\frac{1}{\alpha_{2}}+\frac{1}{\alpha_{3}}\geq 1$, then the only possibilities for the triple $\{\alpha_{1},\alpha_{2},\alpha_{3}\}$ are $\{2,3,3\}$, $\{2,3,5\}$, and $\{3,3,3\}$ since at most one of $\alpha_{i}$ is even because $\Sigma_{2}(K)$ is a $\mathbb{Z}/2\mathbb{Z}$-homology sphere. This case is impossible since neither $T(2,3,3)$, $T(2,3,5)$, nor $T(3,3,3)$ contains an element of order $7$.
When $\frac{1}{\alpha_{1}}+\frac{1}{\alpha_{2}}+\frac{1}{\alpha_{3}}<1$, $\bar{\varphi}(T(3,3,7))=T(\alpha_{1},\alpha_{2},\alpha_{3})\subset\mathit{PSL}(2,\mathbb{R})$ is a hyperbolic triangle group generated by two elliptic elements $a=\bar{\varphi}(y)$ and $b=\bar{\varphi}(z)$ of orders $3$ and $7$, respectively. Then, up to taking suitable powers $u=a^{k}$ and $v=b^{\ell}$ to normalize the matrix representatives of $u$ and $v$ in $\mathit{SL}(2,\mathbb{R})$, at least one of Cases (I)–(VII) in [41, Theorem 2.3] holds. Since the triangle group $T(\alpha_{1},\alpha_{2},\alpha_{3})$ is co-compact, Case (II) is impossible by [41, Figures 2 and 3]. Cases (III) and (VI) do not hold because the generators $u$ and $v$ do not have the same order. Case (IV) is not possible because neither of the generators $u$ nor $v$ has order $2$. Cases (V) and (VII) correspond to the case where $u$ and $v$ generate a hyperbolic triangle group of type $T(2,3,7)$ by [41, Section 3]. Hence, among Cases (II)–(VII), the only possibility left is that $\bar{\varphi}(T(3,3,7))=T(2,3,7)$, which is ruled out by Lemma 7.14 below.
In Case (I), $uv$ has extreme negative trace and by [41, Proposition 2.2], $u$ and $v$ generate a hyperbolic triangle group $T(3,7,n)$. The epimorphism $\bar{\varphi}\colon T(3,3,7)$ $\twoheadrightarrow T(3,7,n),$ implies that $\chi(S^{2}(3,3,7))\geq-\chi(S^{2}(3,7,n))$ by [57, Lemma 2.5]. So, $1-(\frac{1}{3}+\frac{1}{3}+\frac{1}{7})\geq 1-(\frac{1}{3}+\frac{1}{7}+\frac{1}{n})$, and hence $n\leq 3$. Assume $n=3$. Since triangle groups are hopfian, the epimorphism $\bar{\varphi}$ is an isomorphism. Furthermore, the restriction of $\tilde{\varphi}\colon\pi_{1}(\Sigma_{2}(11a_{103}))\twoheadrightarrow\pi_{1}(\Sigma_{2}(K))$ to the center $Z(\pi_{1}(\Sigma_{2}(11a_{103})))\cong\mathbb{Z}$ is injective since $\pi_{1}(\Sigma_{2}(K))$ is torsion-free. Therefore, the epimorphism $\tilde{\varphi}$ is an isomorphism and, by the five lemma, so is $\varphi\colon G^{\mathrm{orb}}(11a_{103})\twoheadrightarrow G^{\mathrm{orb}}(K)$. Therefore, like in Cases (II)–(VII) above, the only possibility is that $\bar{\varphi}(T(3,3,7))=T(2,3,7)$, which is ruled out by Lemma 7.14 below. This completes the proof. ∎
@@
There is no epimorphism $\tilde{\varphi}\colon\pi_{1}(V(0;\frac{9}{7};\frac{2}{3},\frac{1}{3},\frac{2}{7}))\twoheadrightarrow\pi_{1}(V(0;e_{0};\frac{\beta_{1}}{2},\frac{\beta_{2}}{3},\frac{\beta_{3}}{7}))$.
@@
Suppose, to the contrary, that there is an epimorphism $\tilde{\varphi}\colon\pi_{1}(M)\twoheadrightarrow\pi_{1}(N)$. Since $\beta_{2}$ is coprime to $3$, $|H_{1}(N;\mathbb{Z})|=21\beta_{1}+14\beta_{2}+6\beta_{3}$ is coprime to $3$. Moreover, $|H_{1}(N;\mathbb{Z})|$ must divide $|H_{1}(M;\mathbb{Z})|=81$, and hence $|H_{1}(N;\mathbb{Z})|=1$. So $N$ is the Brieskorn homology sphere $V(0;\frac{1}{42};\frac{1}{2},\frac{-1}{3},\frac{-1}{7})$, see [35, Theorem 7.12]. Since $M$ and $N$ are aspherical, the epimorphism $\tilde{\varphi}$ can be realized by a map $f\colon M\to N$. Since both $\pi_{1}(M)$ and $\pi_{1}(N)$ have rank $2$, the $\pi_{1}$-surjective map $f$ is of non-zero degree by [56, Theorem 2.1]. Then, by [58, Theorem 3.2 and Remark 2], $f$ is homotopic to $g\circ\pi$, where $\pi$ is a composition of finitely many vertical pinches and $g$ is a fiber-preserving branched covering. Since $M$ has only three exceptional fibers, a vertical pinch on $M$ would produce a Seifert fibered manifold with $\leq 2$ exceptional fibers, and hence with cyclic fundamental group. Therefore, up to homotopy, $f$ cannot factorize through a vertical pinch since $\pi_{1}(N)$ is not cyclic. It follows that $f$ is homotopic to a fiber-preserving branched covering $g$. Let $d_{1}$ be the degree of $g$ along the fiber and $d_{2}$ the degree of the induced map $\bar{g}\colon S^{2}(3,3,7)\to S^{2}(2,3,3)$. Then, $\frac{4}{21}=|\chi(S^{2}(3,3,7))|\geq d_{2}|\chi(S^{2}(2,3,7))|=\frac{d_{2}}{42}$. This implies that $8\geq d_{2}$. Moreover, $\frac{9}{7}=e_{0}(M)=\frac{d_{2}}{d_{1}}e_{0}(N)=\frac{d_{2}}{42d_{1}}$ by [52, Theorem 1.2]. Then $d_{1}=\frac{d_{2}}{54}\leq\frac{8}{54}$, which is impossible. ∎
Proposition 7.10 allows us to give some lower bounds for the bridge number, genus, and unknotting number of the partial knot of a symmetric union presentation. For a knot $K$, let $b(K)$ denote the bridge number, $g(K)$ the Seifert genus, and $u(K)$ the unknotting number of $K$.
Corollary 7.15.
If $K$ admits a symmetric union presentation with partial knot $K_{D}$, then
  1. (1)
    $b(K_{D})\geq\frac{1}{2}\operatorname{rank}H_{1}(\Sigma_{2}(K);\mathbb{Z})+1$.
  2. (2)
    $g(K_{D})\geq\frac{1}{4}\operatorname{rank}H_{1}(\Sigma_{2}(K);\mathbb{Z})$.
  3. (3)
    $u(K_{D})\geq\frac{1}{2}\operatorname{rank}H_{1}(\Sigma_{2}(K);\mathbb{Z})$.
@@
The proofs follow from Proposition 7.10 and the following well-known facts. First, $b(K_{D})$ verifies the inequality $b(K_{D})\geq\operatorname{rank}\pi_{1}(\Sigma_{2}(K_{D}))+1$, see [11, Lemma 1.6]. Next, for a Seifert matrix $S$ of $K_{D}$, the sum $S+S^{T}$ gives a presentation matrix for the $\mathbb{Z}$-module $H_{1}(\Sigma_{2}(K_{D});\mathbb{Z})$, and therefore $2g(K_{D})\geq\operatorname{rank}H_{1}(\Sigma_{2}(K_{D});\mathbb{Z})$. Finally, Wendt’s inequality [63] shows that $u(K_{D})\geq\operatorname{rank}H_{1}(\Sigma_{2}(K_{D});\mathbb{Z})$ (see also [38]). ∎
Remark 7.16.
In some cases, Wendt’s inequality is powerful. For $K=13a_{1786}$, we have $u(K)\leq 3$ since a single crossing change produces $10_{75}$ whose unknotting number is $2$. According to KnotInfo, the smooth $4$-genus of $K$ equals $1$, and hence one cannot use it to determine $u(K)$. On the other hand, the invariant factors of $H_{1}(\Sigma_{2}(K);\mathbb{Z})$ are $\{3,3,33\}$, and thus Wendt’s inequality implies $u(K)=3$.
8. The bridge number and preorder $\succeq$